OpenStudy (anonymous):

Y = 5000 + 7.5X...anybody have the patience to solve it step by step?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

K, thanks!

OpenStudy (anonymous):

What is it you are trying to solve

OpenStudy (anonymous):

This is a linear equation

OpenStudy (anonymous):

Given above equation I have to I dentify Y intercept, x intercept and slope...do I get these directly from the equation w/o attempting to solve antything?

OpenStudy (anonymous):

I understand

OpenStudy (anonymous):

To find the y-intercept you must make x=0 so Y = 5000 + 7.5X becomes Y = 5000 + 7.5(0) Therefore y = 5000

OpenStudy (anonymous):

To find the x intercept you make y=0 Therefore 0 = 5000 + 7.5X Subtract 5000 from both sides -5000 = 7.5x Divide both sides by 7.5 x = -666.66666.......

OpenStudy (anonymous):

Question: making x= 0...is that a rule or random b/c you know this stuff?

OpenStudy (anonymous):

Well if you want to find where the equation crosses the y-axis (y-intercept) x HAS to be zero. Because if x is any other number, say 2. You wouldn't be on the axis you would be in one of the quadrants. Same with finding the x-intercept.

OpenStudy (anonymous):

Gianfranco and Malevolence19 thank you both! You prevented further burning off of brain cells! :)

OpenStudy (anonymous):

Niente

OpenStudy (anonymous):

Trying to figure out how to thank you guys and say I have a good answer!

OpenStudy (anonymous):

Toni 63 to answer your question it is a logical rule which we always use in this instance. .Thanks for the medal